OpenStudy (anonymous):

xplain, in complete sentences, how you would factor 27x^4-90x^2+75

myininaya (myininaya):

Okay first it easier to look at \[27u^2-90u+75\] and the way we can do that is making the substitution \[u=x^2\] so first we ask ourselves what is the gcf of 27,90,75?

OpenStudy (anonymous):

I will stare at it and keep staring and then write \( 3 \left(3 x^2-5\right)^2 \)

myininaya (myininaya):

so fool just mentioned the gcf which is 3

OpenStudy (mr.math):

stare at it and keep staring! LOL!

OpenStudy (anonymous):

ok

myininaya (myininaya):

\[3(9u^2-30u+25)\] now we look at the inside (the thing in the grouping symbols) \[9u^2-30u+25\] note we will bring the 3 down later ok so a=9,b=-30, c=25 We need to find two factors of a*c that have product a*c and have sum b. \[a \cdot c=9(25)=45(5)=15(15)=-15(-15)\] b=-15+(-15)=-15-15 (whatever!) so Then we replace bu with -15u-15u \[9u^2-30u+25=9u^2-15u-15u+25\] Then we factor by grouping \[3u(3u-5)-5(3u-5)=(3u-5)(3u-5)=(3u-5)^2\] so we have (remember to bring down mr.3 ) \[3(3u-5)^2\] we should probably put this back in terms of x so recall we let \[u=x^2\] so replacing u in \[3(3u-5)^2\] with \[x^2\] we get \[3(3x^2-5)^2\] which is what Fool had

OpenStudy (turingtest):

Bravo myin, determined to explain it :)

myininaya (myininaya):

lol